No of atoms in 28g of nitrogen gas
A.2
B.1
C. 6.022 x 10 raise to power 23
D. 12.044 x 10 raise to power 23
1
Expert's answer
2021-08-06T05:15:19-0400

A mole (mol) is the amount of a substance that contains 6.02 × 10 23 representative particles of that substance.

n = m / M

M (N2) = 28 g/mol

n (N2) = 28 / 28 = 1 mol

Therefore, 28 g of Nitrogen gas contains (C. 6.022 x 10 raise to power 23) atoms.